10        Which of the following atoms can participate in hydrogen bonds?

a.

carbon

b.

oxygen

c.

nitrogen

d.

b and c

e.

all of the above

11        If the pH of a solution is 10 and the concentration of protons is increased 10 fold by the addition of acid, what is the new concentration of protons?

a.

10-11 M

b.

10-9 M

c.

1 M

d.

10-10M

e.

none of the above
Answer #1

10. Hydrogen bond is a dipole-dipole attraction between hydrogen atom attached to an electronegative atom and an electronegative atom.

Carbon participates in covalent bonding with hydrogen as it has four electrons in its outermost shell. Hence, it has low electronegativity. However, the hydrogen in C-H bond can form hydrogen bonds with O, N or F.

Oxygen and nitrogen are highly electronegative. Hence, they can form hydrogen bonds.

Right choice: d. b and c.

11. pH= -log [H]

pH=10

[H]= 10-pH= 10-10 M

Each pH increases of 1, is 10 fold decrease in protons (H+), while a decrease of 1 pH scale, is an 10 fold increase in protons.

If [H] is increases 10-fold, then the pH should decrease by 1 pH unit.

Hence, pH will be 9 in this case.

[H+] = [H]= 10-pH= 10-9 M

Right Choice= b. 10-9 M

10. Hydrogen bond is a dipole-dipole attraction between hydrogen atom attached to an electronegative atom and an electronegative atom.

Carbon participates in covalent bonding with hydrogen as it has four electrons in its outermost shell. Hence, it has low electronegativity. However, the hydrogen in C-H bond can form hydrogen bonds with O, N or F.

Oxygen and nitrogen are highly electronegative. Hence, they can form hydrogen bonds.

Right choice: d. b and c.

11. pH= -log [H]

pH=10

[H]= 10-pH= 10-10 M

Each pH increases of 1, is 10 fold decrease in protons (H+), while a decrease of 1 pH scale, is an 10 fold increase in protons.

If [H] is increases 10-fold, then the pH should decrease by 1 pH unit.

Hence, pH will be 9 in this case.

[H+] = [H]= 10-pH= 10-9 M

Right Choice= b. 10-9 M